a population consists of three numbers (2,5,9). Consider all possible samples of size 2 which can be drawn from the population. Find the variance of the sampling distribution of the sample means.

1
Expert's answer
2021-11-09T11:44:43-0500

There are "\binom{N}{n}=N!/(N-n)!*n!" possible ways of drawing a sample of size "n" from a population size of "N".

In this question, "N=3" and "n=2". Therefore, number of ways to draw samples of size "n=2" is,

"\binom{3}{2}=3!/2!=3"

The samples are,

(2,5),(2,9),(5,9)

The population mean"(\mu)" is given as,

"\mu=\sum(x)/N=(2+5+9)/3=16/3=5.3333"

The population variance"(\sigma^2)" is given as,

"\sigma^2=\displaystyle\sum^2_{i=1}(x_i-\mu)^2/N=24.6667/3=8.2222"

Define the variance of the sampling distribution of the sample means by "\sigma^2_m" then it is given by,

"\sigma^2_m=\sigma^2/n=8.22222/2=4.1111(4\space dp)"

Thus, the variance of the sampling distribution of the sample means is "\sigma^2_m=4.1111"